OpenStudy (ggray):

What is the value of the function y= 0.5x-10 when x is 9? Please help will medal

OpenStudy (rayaz):

Put x=9 into the equation y. y = 0.5(9)-10

OpenStudy (ggray):

Would it be 5.5 or -5.5

Mehek (mehek14):

y = 0.5 * 9 - 10 y = 4.5 - 10 y = -5.5

OpenStudy (ggray):

Thanks!!

Mehek (mehek14):

yw :)